You want to design an oval racetrack such that 3200 lb racecars can round the turns of radius 1000 ft at 1.00 x102 mi/h without the aid of friction. You estimate that when elements like downforce and grip in the tires are considered the cars will round the turns at a maximum of 175 mi/h. Find the banking angle e necessary for the racecars to navigate these turns at 1.00 x102 mi/h and without the aid of friction. Number This banking and radius are very close to the actual turn data at Daytona International Speedway where 3200 lb stock cars travel a the turns at about 175 mih. What additional radial force is necessary to hold the racecar on the track at 175 mi/h? Number
Answer #1

m =3200 lb = 1451.5 kg
100mph = 146.667 ft/s
Since there is no friction,

tan theta = centripetal acceleration/g

theta = arctan(a/g) = arctan(v^2/rg)

= arctan(146.667^2/(1000*32))

= 33.91 degree
B] 175 mph = 78.232 m/s , 1 ft = 0.3048 m

now tan theta = a/g

a = g* tan 33.91 degree

ma = mg* tan 33.91 degree

= 3200 * tan 33.91degree = 2151 lb = 975.7*9.8 N = 9562 N

Centripetal force = mv^2/r = 1451.5*78.232^2/(1000*0.3048) = 29145 N

additional force required = 29145 - 9562 N = 19583 N answer